Gastric acid pH can range from 1 to 4, and most of the acid is HCl.
For a sample of stomach acid that is 1.42×10−2 M in HCl, how many moles of HCl are in 18.3 mL of the stomach acid?
Answer #1

M(HCl) = 1.42*10^-2 M

V(HCl) = 18.3 mL = 0.0183 L

Use:

mol of HCl = M(HCl)*V(HCl)

= 1.42*10^-2 M * 0.0183 L

= 2.60*10^-4 mol

Answer: 2.60*10^-4 mol